OpenStudy (anonymous):

Select the coordinates of two points on the line y = −2. (2, −2) and (−2, 2) (−2, −2) and (−2, 0) (2, −2) and (0, −2) (−2, 2) and (−2, −2)

OpenStudy (anonymous):

the form is (x,y) so it the third one

OpenStudy (anonymous):

@ArifL thank you!! i understand now!

jimthompson5910 (jim_thompson5910):

every point on the line y = -2 has a y coordinate of -2 some examples are (5,-2), (10,-2), (144,-2), etc. The x coordinate can be anything you want. The y coordinate must be -2

OpenStudy (dayakar):

y= -2 shows a strightline parallel to X-axis passing through \[\left( x ,-2\right)\] therfore answer is \[\left( 2,-2 \right)and \left( 0,-2 \right)\]

OpenStudy (anonymous):

Every point where y=-2 so its the third one, (2, −2) and (0, −2)

OpenStudy (anonymous):